# A compact model of Escherichia coli core and biosynthetic metabolism

### v1

@@ A.5.2 Adjusting turnover numbers for the reference set of flux distributions
A.5.3 Conversion to apparent turnover numbers
The above procedure produces (after conversion back to a linear scale) a set of adjusted turnover numbers, $\mathbf{k}_{\rm cat}^{*}$ and scalings, $\boldsymbol{\sigma}^{*}$. In order to facilitate the physical interpretation of the resulting parameter set, we express the fitted scalings for the $j$th condition, $\sigma_{j}^{*}$ as the product of two terms:
$\sigma_{j}^{*}\equiv\bar{\sigma}^{*}~{}\hat{\sigma}_{j}^{*}$ (30)
Here, $\bar{\sigma}^{*}$ is the geometric mean of the scalings across conditions, which we interpret as typical enzyme saturation level, while $\hat{\sigma}_{j}^{*}$ is a residual scaling factor fluctuating around $1$ which required to account for differences in total measured enzymes across conditions. Based on this reparametrisation of the scaling, we now define a set of adjusted apparent turnover numbers, $\mathbf{k}_{\rm cat}^{\rm app}$, which now incorporate saturation information, as:
$\mathbf{k}_{\rm cat}^{\rm app~{}*}\equiv\bar{\sigma}^{*}~{}\mathbf{k}_{\rm cat}^{*}$ (31)
A.5.4 Potential extensions
We conclude this section by noting that procedure described above assumes that the original parameter set is sufficiently ”good” to produce a realistic flux distribution to use as a reference for the adjustment step. If this is not the case, then a multi-start approach can be implemented, where multiple turnover parameter sets are first generated by perturbing the original parameter set, and each is used to generate a separate set of flux distributions. Each reference set is then provided as an input to problem (29), and the solution achieving the lowest objective chosen in the end. The perturbed parameter set may be generated either randomly (for example, by adding log-normal noise to the original turnover parameter vector) or systematically. The latter could be achieved, for example, by identifying reactions with zero-predicted flux but high measured abundance of the associated enzyme. By systematically increasing the corresponding turnover number, one can ”encourage” these reactions to be included in the reference set, potentially leading to the exploration of more relevant reference set than achieved by random perturbation.
Note that, for the purposes of this work, we limited ourselves to the original parameter set and thus did not explore this potential heuristic.

### v2

@@ A.5.4 Estimating typical enzyme efficiencies for the reference set of flux distributions
A.5.5 Conversion to apparent turnover numbers
The above procedure produces (after conversion back to a linear scale) a set of adjusted turnover parameters, $\mathbf{k}_{i}^{*}$ and scalings, $\boldsymbol{\theta}^{*}$. In order to retrieve the typical enzyme efficiencies and condition-specific scaling factors introduced in 25, we simply parametrise the solution by factorising the scaling terms $\theta_{j}$ as
$\sigma_{j}^{*}\equiv\bar{\sigma}^{*}~{}\tau_{j}^{*}$ (36)
Here, $\bar{\sigma}^{*}$ is the geometric mean of the scalings across conditions, which we interpret as typical enzyme saturation level across conditions, while $\tau_{j}^{*}$ is a residual scaling factor fluctuating around $1$ which required to account for differences in total measured enzymes across conditions. The typical enzyme efficiencies ($\kappa_{i}$) introduced in 25 are then simply recovered by incorporating the $\bar{\sigma}^{*}$ constant into the fitted turnover parameters:
$\kappa_{i}\equiv\bar{\sigma}^{*}~{}k_{i}^{*}$ (37)
A.5.6 Potential extensions
We conclude this section by noting that the procedure described above assumes that the original parameter set – the set of $k_{\rm cat}$ values as proxies for apparent $k_{\rm cat}$ values in section A.5.3 – is sufficiently good to produce a realistic flux distribution to use as a reference for the adjustment step. If this is not the case, then a multi-start approach can be implemented, where multiple turnover parameter sets are first generated by perturbing the original parameter set, and each is used to generate a separate set of flux distributions. Each reference set is then provided as an input to problem (35), and the solution achieving the lowest objective is chosen in the end. The perturbed parameter set may be generated either randomly (for example, by introducing log-normal noise in the original turnover parameter vector) or systematically. The latter could be achieved, for example, by identifying reactions with zero-predicted flux but high measured abundance of the associated enzyme. By systematically increasing the corresponding turnover number, one can ”encourage” these reactions to be included in the reference set, potentially leading to the exploration of more relevant reference distributions than achieved by random perturbation.
Note that, for in this work, we limited ourselves to the original parameter set and thus did not explore this potential heuristic.
@@ Appendix C Supplementary Tables
[…]
$\rightarrow$ RED-THIOREDOXIN-MONOMER
$\rightarrow$ RED-THIOREDOXIN2-MONOMER